40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Apr 10 19:14:08 2011, Koji, Update, Green Locking, New Green PD test1 SGD444A_240.pngSGD444A_240_test.png
    Reply  Mon Apr 11 09:49:53 2011, rana, Update, Electronics, New Green PD test1 
       Reply  Mon Apr 11 11:34:05 2011, Koji, Update, Electronics, New Green PD test1 SGD444A_240_test.png
          Reply  Mon Apr 11 21:13:15 2011, Koji, Update, Electronics, New Green PD test1 
Message ID: 4506     Entry time: Sun Apr 10 19:14:08 2011     Reply to this: 4507
Author: Koji 
Type: Update 
Category: Green Locking 
Subject: New Green PD test1 

I started to modify another green PD set.

It so far has the transimpedance of 240 Ohm on CLC409 for the RF output.

It shows the BB output upto ~100MHz.
The measurement shows the transimpedenca of ~90Ohm which is ~25% smaller than the expected gain of 120Ohm.
It is calibrated based on the transimpedances of Newfocus 1611 (10kOhm and 700Ohm for AF and RF).

The next step is to change the transimpedance resister to 2k and replace the PD to S3399 Si PD, which has the diameter of 3mm.
Then, the noise level will be measured. (and replace the RF opamp if necessary)

 

Attachment 1: SGD444A_240.png  280 kB  | Hide | Hide all
SGD444A_240.png
Attachment 2: SGD444A_240_test.png  40 kB  | Hide | Hide all
SGD444A_240_test.png
ELOG V3.1.3-